Next meeting chair: Mas
Attended: Lim, Jay, Jamie, Mas, Prad, Simin
- Use new wiki for meetings. Suggestion: prepend new entries
- Ho26m receiver updates? Should be ready for AOV next week. Seems working in the lab.
- Flash end of support, ptz therefore breaks.
- Hobart12 samplers bad - removed Hb dbbc3 to allow cold air flow through (which should help).
- Katherine12 samplers bad - data still useable. Need to work out better way to detect the false. Clock break happens. Soft reset of dbbc3ke.reset.py script instead of reconfiguring. Power level looks quite marginal. Nothing the observers can do for the time being yet.
- Yarragadee network out and pointing issue. Pointing models corrected and fine.
- Hobart12 - 1145UT Issues with antenna, system had to be rebooted. (Eloise, done by Warren) - Switch overload? (no issue found for now, Jamie will investigate the cabling further)
- Katherine12 - No data was being recorded. Multiple resets of dbbc and running setup seemed to work. Recording started at 1954UT - probably with flexbuff and datastream, changing from spiral to aum mode could be the issue. Happens occasionally. Easiest thing to do could be just kill jive5ab and restart by StartJ5.
- Hobart 26 - The receiver was changed to the L-band an hour prior to the experiment. So the timed start recorded at the wrong frequency for the first hour. Some monitoring related issues showed up. Experiment started recording at 1922UT. – No receivers so we shouldn't have observed.
- Katherine12 - antenna stuck briefly before experiment started - had to reboot system (antenna=open didn't work) (Eloise)
- Katherine12 - 2220UT Antenna communication issues since early morning, resolved with antenna=open (Eloise)